OpenStudy (anonymous):

Can someone check my work? What is the tangent line to the curve at the given point. y=sqrt(x) (1,1) My answer is... f(x)-f(a)/x-a sqrt(x)-1/1-1=sqrt(x)-1/0 is this right? If not, can you correct me?

OpenStudy (anonymous):

I don't think that's quite it... looking at it now though

OpenStudy (anonymous):

differentiate it \[\frac{dy}{dx}=\frac{1}{2\sqrt{x}}\] then sub in \(1\)

OpenStudy (anonymous):

yes, thanks! my equation typing energy was running low :)

OpenStudy (anonymous):

1/2? is it y=1/2x+3/2?

OpenStudy (anonymous):

y=1/2x-3/2 correction

OpenStudy (anonymous):

the derivative of the curve gives the slope of the curve

OpenStudy (anonymous):

I know that.. which means the slope is 1/2

OpenStudy (anonymous):

so you want the slope of the curve f(x) at point (1,1), so you put x = 1 into the dy/dx expression and get 1/2

OpenStudy (anonymous):

that's the slope. Then use that same point to build the y = mx + b line equation for the tangent

OpenStudy (anonymous):

y=1/2x+1/2?

OpenStudy (anonymous):

yeah I think that's better... good fix.

OpenStudy (anonymous):

thanks.

OpenStudy (anonymous):

glad to help :)